OpenStudy (anonymous):

Algebra 2 please help

OpenStudy (anonymous):

With what?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

kk.

OpenStudy (anonymous):

simplify\[3(\log_{4}9+\log_{4}t) \]

OpenStudy (anonymous):

sorry don't know.

OpenStudy (dumbcow):

\[\log a + \log b = \log(ab)\] and \[n*\log(x) = \log(x^{n})\]

OpenStudy (anonymous):

\[\log_{4}(9t)^{3} \]would it be

OpenStudy (dumbcow):

yep you got it :)

OpenStudy (anonymous):

thank you :)